What is the strength​ (expressed as a ratio in lowest​ terms) of a 300 milliliter solution that contains 150 grams of sodium​ chloride?
Answer #1

Strength will be expressed as w/v ratio.

w/v = weight of solute/ solution volume

w/v = 150/300 ml = 1/2 w/v

Strength is 1:2 w/v.